OpenStudy (anonymous):

Write using exponents. (–4)(–4)

OpenStudy (anonymous):

@Tom_Boy_Rebel I must show work. HELP!!!!

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

Let me see.

OpenStudy (anonymous):

=.)...yay

OpenStudy (anonymous):

Okay. If you had a problem that asked to write (-5)(-5)(-5)(-5) as a exponent, then your answer would be -5^how many (-5) there are...So its -5^4.

OpenStudy (anonymous):

Oh he is teaching it sorry

OpenStudy (anonymous):

So, \[x \times x = x^2\] \[3 \times 3 = 3^2\] So I believe if you do: \[(-4) \times (-4) = (-4)^2\] That should be the answer :)

OpenStudy (anonymous):

Hope this helped! Have a great day! Also a medal would be much appreciated! Just click best response next to my answer. Thank You! @ciaramahogany

OpenStudy (anonymous):

I am a girl but thats okay :) @ZacharySanders Also thank you for helping. If someone is answering that doesn't mean you can't help as well :)

OpenStudy (anonymous):

Oh so sorry I didn't read your name fully

OpenStudy (anonymous):

Lol its fine.

OpenStudy (anonymous):

Thank you so much @tom_boy_rebel

OpenStudy (anonymous):

No Problem! :) @ciaramahogany

OpenStudy (anonymous):

@Tom_Boy_Rebel ....does this look right? (–4)(–4) x*x=x>2 3*3=3>2 (-4)*(4)=(-4)>2